# A Conflicts-free, Speed-lossless KAN-based Reinforcement Learning Decision System for Interactive Driving in Roundabouts

### v1

@@ IV-C4 Update Rule
After each EV’s action decision, the action with the next highest priority is selected. This process of actions inspector repeats until EV have safely navigated through the roundabout. The actions inspector helps the EV avoid conflicts by continuously replacing dangerous actions with safer alternatives, as summarized in the Algorithm 2.
In summary, the proposed route planner and action inspector provide a comprehensive solution for safe and efficient navigation of autonomous vehicles in complex roundabout scenarios. The route planner’s initial lane selection based on TTC ensures a safe entry into the roundabout, while the modified BFS algorithm optimizes the path by considering both distance and traffic density. The lane selection mechanism, which incorporates traffic density and lane change costs, enables strategic and safe lane changes within the roundabout. The action inspector continuously monitors the safety of planned actions and replaces dangerous maneuvers with safer alternatives, minimizing the risk of collisions. By integrating global route planning with localized lane change decisions based on real-time traffic data, the proposed system demonstrates exceptional adaptability to varying traffic conditions.
V MPC for enhancing DRL Performance
This section introduces the robust control for AVs including the vehicle dynamic model and MPC. The MPC controller considers the vehicle dynamics, collision avoidance, and other constraints in its optimization process. It predicts the future states of the EV and surrounding vehicles using the vehicle dynamic model and the actions of neighboring vehicles predicted by the DRL agent. The combination of DRL and MPC in the proposed framework brings several benefits: it allows the DRL agent to focus on high-level decisions while the MPC manages low-level controls; the MPC can correct any imperfections in the DRL agent’s decisions to ensure safe and feasible actions; and MPC provides a reliable, interpretable control strategy based on clear vehicle dynamics and constraints.
The state of the EV at time step $t$ is updated using
$p_{\textrm{EV}}(t+1)$ $=p_{\textrm{EV}}(t)+v_{\textrm{EV}}(t)\cdot\cos(h_{\textrm{EV}}(t))\cdot\Delta t$
$v_{\textrm{EV}}(t+1)$ $=v_{\textrm{EV}}(t)+u(t)\cdot\Delta t$
$h_{\textrm{EV}}(t+1)$ $=h_{\textrm{EV}}(t)+\frac{v_{\textrm{EV}}(t)}{L}\cdot\tan(\delta(t))\cdot\Delta t$ (41)
$v_{\min}$ $\leq v_{\textrm{EV}}(t)\leq v_{\max},~{}a_{\textrm{min}}\leq u(t)\leq a_{\textrm{max}},$
$\delta_{\textrm{min}}$ $\leq\delta(t)\leq\delta_{\textrm{max}},~{}\lVert p_{\textrm{EV}}(t)-p_{\textrm{NV}}(t)\rVert\geq D_{safe}.$ (42)
$\hskip 50.00008pt\min J_{c}$
$\mathbf{s.t.}~{}$ (41)$,0<v_{\textrm{AV}}(k)\leq v_{\textrm{max}},a_{\textrm{min}}\leq u(k)\leq a_{\textrm{max}},$
$\quad\delta_{\textrm{min}}\leq\delta(k)\leq\delta_{\textrm{max}},NV\in N_{V},k\in[0,N_{p}-1]$ (43)
with the cost function $J_{c}=\sum_{k=0}^{N_{p}-1}(v_{\textrm{AV}}(k)-v_{\textrm{AV}}^{*})^{2}+\sum_{k=0}^{N_{p}-1}\sum_{i=1}^{N_{v}}(\lVert p_{\textrm{SV}}^{i}(k)-p_{\textrm{AV}}^{i}(k)\rVert-D_{safe})^{2}+\lambda\sum_{k=0}^{N_{c}-1}u^{2}(k)$. $N_{p}$ and $N_{c}$ represent the prediction horizon and control horizon, respectively. $N_{p}$ is the prediction horizon and $N_{c}$ is the control horizon. $v_{\textrm{AV}}^{*}(k)$ is the target speed, $N_{v}$ is the number of NVs, and $\lambda$ is a weighting factor for the acceleration penalty. In our experiments, we set $N_{p}$ as 10 and the $N_{c}$ as 5. The entire control process is summarized in Algorithm 3.
@@ V MPC for enhancing DRL Performance
[…]
@@
[…]
@@
18:         $J_{c}\leftarrow J_{c}+(\lVert p_{\textrm{NV}}^{i}(k)-p_{\textrm{EV}}^{i}(k)\rVert-D_{safe})^{2}$
@@
[…]
27:else
[…]
29:end if
@@ VI Simulation Results
In this section, we assess the performance of the proposed K-DQN algorithm by evaluating its training efficiency and collision rate within the roundabout driving scenario, within the roundabout driving scenario depicted in Section II. The inner roundabout’s road radius is 40 meters, and the outer roundabout’s road radius is 48 meters, with both lanes having a width of 4 meters. Vehicles that exit the roundabout will be removed from the AV’s view, although their kinematics will continue to be updated. Specifically, we examine three situations with different sectional configurations and varying levels of traffic density, characterized by distinct safety and control mechanisms, and different numbers of initial vehicles. The situations are defined as follows:
  • Functional mechanism validation on hard mode: The proposed K-DQN is compared with K-DQN without a safety inspector and without the MPC control module.
  • Normal mode validation: The proposed K-DQN is compared with benchmark algorithms with seven initial vehicles in the roundabout as depicted in Fig. 5(a).
  • Hard mode validation: The proposed K-DQN is compared with benchmark algorithms with eleven initial vehicles in the roundabout as depicted in Fig. 5(b).
@@
Figure 5: Validation settings. (a) is the normal mode settings for the roundabout scenario (six initial HDVs); (b) is the hard mode settings (ten initial HDVs).
@@
The benchmarks used for comparison in the normal and hard mode validations include PPO [21], A2C [31], ACKTR [32], and DQN [33]. The performance metrics used for evaluation include training convergence rate, collision rate, average speed, and reward values during training and evaluation. Considering the inherent risks associated with real-world vehicles and the constraints imposed by legal regulations, scenario-based virtual testing offers significant benefits, such as, precise environmental replication and enhanced testing efficiency. Therefore, this study employs scenarios developed on the Highway virtual simulation platform [34]. At the end of each episode, the vehicles and their velocities are slightly randomized at their spawn points to enhance the generalization capability of our proposed model. The computer and environment setup for this study include Python 3.6, PyTorch 1.10.0, Ubuntu 20.04.6 LTS OS, a 12th generation 16-thread Intel® Core™ i5-12600KF CPU, an NVIDIA GeForce RTX 3090 GPU, and 64GB of RAM.
VI-A Functional Mechanism Validation
@@
This section describes experiments to evaluate the crucial functions of the safety inspector and MPC of the proposed system in hard mode. We divide the experiments into four validations: Validation 1 evaluates training convergence of K-DQN, K-DQN without safety inspector, and K-DQN without MPC. Validation 2 tests the stability of the speed variations. Validation 3 compares the reward across the evaluation. Validation 4 analyzes the number of collisions.
Figure 6: Curves of reward during training of K-DQN, and K-DQN with no MPC and no inspector, respectively. The shadow region denotes the standard deviations over 3 random seeds. The curve are smoothed over the last 9 evaluation episodes.
Figure 7: Curves of speed during training of K-DQN, and K-DQN with no MPC and no inspector, respectively.
Figure 8: Curves of reward during evaluation of K-DQN, and K-DQN with no MPC and no inspector, respectively.
@@
Figure 9: Comparison of our performances with SOTA algorithms during 10000 training episodes on normal mode scenario. (a) comparison reward ascending during training (b) comparison speed change rate during training, and (c) comparison collision rate change during training.
Validation 1: Training convergence. To better evaluate performance in the case where surrounding HDVs exhibit random driving behavior, we conducted tests using three random seeds and varied the generated scenarios. Fig. 6 illustrates the comparison of training curves between the proposed K-DQN, the K-DQN without a safety inspector, and the K-DQN without MPC. As expected, the proposed K-DQN design outperforms both the K-DQN without a safety inspector and the K-DQN without MPC, delivering higher peak rewards and faster convergence. The training curves of the proposed K-DQN exhibit smooth convergence and low variance across different random seeds, indicating a stable training process. In contrast, the K-DQN without a safety inspector and the K-DQN without MPC show more fluctuations in their training curves and slower convergence rates. It is evident that the absence of a safety inspector leads to the least satisfactory training outcomes. The safety inspector effectively reduces collisions, aligning with the observation that the loss due to collisions is greater than the loss of efficiency during training. The proposed K-DQN surpasses the peak reward of the K-DQN without MPC in approximately 4000 epochs, indicating its superior training performance.
Validation 2: Stability of the speeds variation. Figure 7 illustrates the comparison of speed variation between the proposed K-DQN, the K-DQN without a safety inspector, and the K-DQN without MPC. As expected, the proposed K-DQN design demonstrates more stable speed control compared to the other two networks. It is evident that the absence of a safety inspector leads to the least stable speed variation, resulting in unstable driving. The safety inspector effectively reduces collisions, which aligns with the observation that collisions result in zero-speed events, significantly decreasing the average speed. The K-DQN without MPC exhibits a speed variation around 5 m/s. In contrast, the proposed K-DQN limits the variation to within 2 m/s, enhancing stability.
@@
Method No MPC No Inspector Ours
Collision Rate (%) 9 11 1
@@
Validation 3: Reward values among evaluation. Figure 8 illustrates the comparison of rewards between the proposed K-DQN, the K-DQN without a safety inspector, and the K-DQN without MPC. As expected, the proposed K-DQN design demonstrates significantly higher rewards compared to both the K-DQN without a safety inspector and the K-DQN without MPC. The absence of a safety inspector results in the lowest rewards (below 100) because the AV collides with HDVs when it interacts with environment during the training episodes without actions inspector to replace dangerous actions. The K-DQN without MPC increases the reward to around 125 compared to the K-DQN without a safety inspector. However, the proposed K-DQN achieves rewards around 175 in most sections throughout the evaluation episodes.
Figure 10: Comparison of our performances with SOTA algorithms during 10000 training episodes on hard mode scenario. (a) comparison reward ascending during training, (b) comparison speed change rate during training, and (c) comparison collision rate change during training.
Validation 4: Collision rate. Table II illustrates the number of collisions and the average speed comparison between the proposed K-DQN, the K-DQN without a safety inspector, and the K-DQN without MPC. As expected, the proposed K-DQN design demonstrates significantly fewer collisions while maintaining a relatively high average speed compared to both the K-DQN without a safety inspector and the K-DQN without MPC. It is evident that the absence of a safety inspector results in the lowest average speed and the highest collision rate. The K-DQN without MPC increases the average speed to 22.88 m/s compared to the K-DQN with a safety inspector, which has an average speed of 16.23 m/s. However, the collision rates for the K-DQN without a safety inspector and the K-DQN without MPC are nearly the same, at 0.11 and 0.09, respectively. The proposed K-DQN achieves a collision rate of 0.01, which is significantly lower than the other two comparison algorithms.
@@ VI-B Normal Mode Validation
In this section, we conduct experiments to evaluate the safety and efficiency of the proposed system in normal mode compared to other benchmark DRL algorithms, PPO [21], A2C [31], ACKTR [32], and DQN [33]. We divide the experiments into two validations: Validation 1 evaluates fast-convergence training, stable and high-speed variation, and collision rate during training. Validation 2 compares the reward levels during evaluation.
Validation 1: Better training performance. To evaluate the training performance of the proposed K-DQN compared to other benchmark algorithms, we conducted tests using three random seeds and varied the generated scenarios. Fig. 9 illustrates the comparison of training curves between the proposed K-DQN and other benchmark algorithms. Fig. 9(a) shows the comparison of rewards during training, where the proposed K-DQN design outperforms other benchmark algorithms, achieving much higher peak rewards and faster convergence. The DQN has the poorest performance, with rewards around 80. The A2C and ACKTR exhibit similar reward levels, while the PPO reaches approximately 125, the highest reward among the other benchmark algorithms. The proposed K-DQN surpasses all, achieving rewards over 200. Fig. 9(b) depicts the comparison of speed variation during training. The proposed K-DQN design excels, maintaining higher speeds and stable speed variation at the end of training. The PPO demonstrates both the lowest speed and the most unstable speed variation. The A2C, ACKTR, and DQN have similar curves, with the A2C reaching a higher speed of around 18 m/s by the end of 10,000 training episodes. The proposed K-DQN maintains a speed level of approximately 22 m/s at the end, significantly faster than the others. Fig. 9(c) shows the collision rate during training, where the proposed K-DQN design again outperforms the benchmark algorithms by achieving a much lower collision rate. The DQN has the highest collision rate at around 0.55. The ACKTR reduces the collision rate to around 0.35. The A2C and PPO have similar collision rates at around 0.2, with the PPO exhibiting greater variability. The proposed K-DQN maintains a collision rate below 0.05, significantly lower than the others.
Validation 2: Reward values among evaluation. Figure 11 illustrates the reward comparison between the proposed K-DQN and other benchmark algorithms. The proposed K-DQN design demonstrates significantly higher rewards compared to all other benchmark algorithms. The DQN has the lowest reward during the evaluation, falling below 75. The A2C and ACKTR are similar, both increasing the reward to around 75. The PPO has a relatively higher reward of around 100, peaking at 125. However, the reward of the proposed K-DQN fluctuates around 175, significantly surpassing other benchmark algorithms.
Figure 11: Curves of reward during evaluation of K-DQN, and benchmark algorithms on normal mode scenario.

### v2

@@ IV-C Action Inspector
4) Update Rule: After each EV action, the next highest-priority action is chosen. This repeats until the EV safely exits the roundabout, with the action inspector continuously replacing risky actions with safer ones (see Algorithm 2).
The action inspector adapts by predicting $T_{n}$ steps ahead and continuously monitoring trajectories. For dynamic traffic density changes, the inspector updates its safety assessments at each time step using real-time traffic information. The system quickly handles unexpected HDV behaviors using priority-based action replacement and IDM following, which adapts to sudden speed changes of preceding vehicles.
@@
9:for $k\leftarrow 0$ to N1N-1 do
10:  for vehicle in $\mathit{NV}$ do
11:   $\text{action}\leftarrow\text{use\_K-DQN}$
12:   $\text{\_to\_predict\_vehicle\_action}(\text{vehicle})$
@@
14:  $\delta(k)\leftarrow\text{compute\_steering}(\mathit{EV},\mathit{NV})$
15:  $\mathit{EV}.\text{update}(u(k),\delta(k))$
@@
18:   $J_{c}\leftarrow J_{c}+(\|p_{\textrm{NV}}^{i}(k)-p_{\textrm{EV}}^{i}(k)\|_{2}-D_{safe})^{2}$
@@
25:if $\mathit{solution}$ found then
26:  return $\mathit{solution}.\text{value}(u[0])$
27:else
28:  return $\text{PID\_controller}(v_{\textrm{EV}}(0),v_{\textrm{EV}}^{*})$
29:end if
In summary, the proposed system combines route planning and action inspection for safe, efficient AV navigation in roundabouts. TTC-based lane selection ensures safe entry, the modified BFS optimizes paths using distance and traffic data, and lane changes are guided by traffic density and cost. The action inspector monitors safety in real time, replacing risky actions to prevent collisions. By integrating global route planning with localized real-time traffic data-based lane change decisions, the proposed system demonstrates exceptional adaptability to varying traffic conditions.
V MPC for enhancing DRL Performance
This section introduces the robust control for AVs including the vehicle dynamic model and MPC. The MPC controller considers the vehicle dynamics, collision avoidance, and other constraints in its optimization process. It predicts the future states of the EV and surrounding vehicles using the vehicle dynamic model and the actions of neighboring vehicles predicted by the DRL agent. The combination of DRL and MPC in the proposed framework brings several benefits: it allows DRL to focus on high-level decisions while MPC manages low-level controls; MPC can correct any imperfections in DRL decisions to ensure safe and feasible actions; and MPC provides a reliable, interpretable control strategy based on clear vehicle dynamics and constraints [36].
The EV’s state is updated by
$p_{\textrm{EV}}(t+1)$ $=p_{\textrm{EV}}(t)+v_{\textrm{EV}}(t)\cdot\cos(h_{\textrm{EV}}(t))\cdot\Delta t$ (42)
$v_{\textrm{EV}}(t+1)$ $=v_{\textrm{EV}}(t)+u(t)\cdot\Delta t$
$h_{\textrm{EV}}(t+1)$ $=h_{\textrm{EV}}(t)+v_{\textrm{EV}}(t)\cdot\tan(\delta(t))\cdot\Delta t/L$
$v_{\min}$ $\leq v_{\textrm{EV}}(t)\leq v_{\max},~a_{\textrm{min}}\leq u(t)\leq a_{\textrm{max}},$ (43)
$\delta_{\textrm{min}}$ $\leq\delta(t)\leq\delta_{\textrm{max}},~\|p_{\textrm{EV}}(t)-p_{\textrm{NV}}(t)\|_{2}\geq D_{safe}.$
$\hskip 50.00008pt\min J_{c}$ (44)
@@ V MPC for enhancing DRL Performance
with the cost function $J_{c}=\sum_{k=0}^{N_{p}-1}(v_{\textrm{AV}}(k)-v_{\textrm{AV}}^{*})^{2}+\sum_{k=0}^{N_{p}-1}\sum_{i=1}^{N_{v}}(\lVert p_{\textrm{SV}}^{i}(k)-p_{\textrm{AV}}^{i}(k)\rVert-D_{safe})^{2}+\lambda\sum_{k=0}^{N_{c}-1}u^{2}(k)$. $N_{p}$ and $N_{c}$ represent the prediction horizon and control horizon, respectively. $v_{\textrm{AV}}^{*}(k)$ is the target speed, and $\lambda$ is a given weighting factor. In our experiments, we set $N_{p}$ as 10 and the $N_{c}$ as 5. The entire control process is summarized in Algorithm 3.
@@ VI Simulation Results
Figure 5: Validation settings: (a) Normal mode with six initial HDVs; (b) Hard mode with ten HDVs.
c
Figure 6: Performance comparison with different components and prediction steps. (a) reward, (b) speed, and (c) collision rates under different prediction steps ($T_{n}$). The shaded regions denote the standard deviations over 3 random seeds.
All curves are smoothed over the last 9 evaluation episodes.
We evaluate K-DQN in the roundabout scenario described in Section II, focusing on training efficiency and collision rate. All experiments are conducted using three random seeds, with mean results plotted and standard deviations shown as shaded areas. The roundabout follows standard traffic engineering design, with an inner radius of $20\,\mathrm{m}$, outer radius of $28\,\mathrm{m}$, and a lane width of $4\,\mathrm{m}$. All vehicles are modeled as typical passenger cars ($4.7\,\mathrm{m}$ long, $2.1\,\mathrm{m}$ wide), and these dimensions are considered in collision detection and safety distance calculations. The algorithm employs a relative state representation and distance-based safety mechanism, ensuring robustness to varying lane widths. Vehicles exiting the roundabout leave the AV’s observation range but continue updating their kinematics. We examine three scenarios as follows:
  • Ablation study in hard mode: The proposed K-DQN is compared with K-DQN without the action inspector, K-DQN without MPC, and the baseline DQN.
  • Normal mode validation: The proposed K-DQN is compared with benchmarks with seven initial vehicles in the roundabout as depicted in Fig. 5(a).
  • Hard mode validation: The proposed K-DQN is compared with benchmarks with eleven initial vehicles in the roundabout as depicted in Fig. 5(b).
@@
The benchmarks used for comparison in the normal and hard mode validations include PPO [23], A2C [37], ACKTR [38], and DQN [28]. The performance metrics used for evaluation include training convergence rate, collision rate, average speed, and reward values during training and evaluation. Considering the inherent risks associated with real-world vehicles and the constraints imposed by legal regulations, scenario-based virtual testing offers significant benefits like precise environmental replication and enhanced testing efficiency. Therefore, this study employs scenarios developed on the Highway virtual simulation platform [39]. To ensure robust performance under varying traffic conditions, HDV speeds are initialized following a normal distribution around a mean speed of $20\,\mathrm{m/s}$ with a $\pm$15% variation. Similarly, HDV’s positions are randomized with a normal distribution to create diverse traffic patterns. At the end of each episode, the vehicles and their velocities are slightly randomized at their spawn points to enhance the generalization capability of our proposed model. This randomization helps evaluate the model’s adaptability to different traffic speeds and densities, which is crucial for real-world deployment. The computer and environment setup for this study include Python 3.6, PyTorch 1.10.0, Ubuntu 20.04.6 LTS OS, a Intel® Core™ i5-12600KF CPU, an NVIDIA GeForce RTX 3090 GPU, and 64GB of RAM.
VI-A Ablation Study in Hard Mode
@@
This section describes experiments to evaluate the crucial functions of the action inspector and MPC of the proposed system in hard mode. To comprehensively assess the contribution of each component, we include a baseline DQN (using traditional MLP architecture) for comparison with our K-DQN variants. We divide the experiments into four validations: Validation 1 evaluates training performance of K-DQN, K-DQN without action inspector, K-DQN without MPC, and baseline DQN. Validation 2 tests the stability of the speed variations. Validation 3 compares the reward across the evaluation. Validation 4 analyzes the number of collisions.
Figure 7: Rewards of different K-DQN schemes and DQN.
@@
Figure 8: Performance comparison with benchmarks in normal mode: (a) reward, (b) speed, and (c) collision rate.
Validation 1: Training performance. To better assess performance under random HDV behavior, we conducted tests using three random seeds and varied scenarios. Figure 6c(a) compares training curves of the proposed K-DQN with three variants: K-DQN without action inspector, K-DQN without MPC, and baseline DQN with a standard MLP architecture. As expected, the full K-DQN consistently achieves higher peak rewards and faster, more stable convergence. Compared to baseline DQN, the K-DQN without inspector already benefits from the KAN architecture, yielding 10–15% higher rewards. However, the largest gain comes from the action inspector, which filters unsafe actions and reduces collision penalties during exploration. The proposed K-DQN shows low variance across seeds, indicating stable training, while K-DQN without the inspector or MPC suffers from more fluctuations and slower convergence. The baseline DQN performs the worst, confirming that both the KAN architecture and action inspector are essential for robust and efficient learning.
@@
Metrics No MPC No Inspector Ours
Collision Rate (%) 9 11 2
@@
[…]
Figure 6c(c) shows how prediction horizon ($T_{n}$) affects the action inspector. With $T_{n}=8$, the collision rate is lowest and most stable ($\sim 0.02$), allowing ample time for action replacement. Shorter horizons ($T_{n}=6$ and 3) raise the rate to $\sim 0.03$ and $\sim 0.07$, while $T_{n}=0$ performs worst ($\sim 0.11$), highlighting the importance of predictive action inspection.
Validation 3: Reward values among evaluation. Figure 7 compares rewards for K-DQN, its ablated variants, and baseline DQN. K-DQN achieves the highest average ($\sim 175$), showing the benefit of integrating the action inspector and MPC. Removing MPC drops rewards to $\sim 125$, and without the inspector, rewards fall below 100 due to frequent collisions. Lastly, the baseline DQN performs the worst, with rewards mostly staying below 75, underscoring the limitations of naive reinforcement learning without structured safety or strategic reasoning. These results collectively highlight the necessity of combining multi-level safety and planning mechanisms for robust autonomous decision-making.
Figure 9: Rewards for K-DQN and benchmarks: normal mode.
Validation 4: Collision rate. Table II compares collisions and average speed for the proposed K-DQN, K-DQN without the action inspector, and K-DQN without MPC. The full K-DQN achieves the lowest collision rate at $2\%$ with a solid average speed of $16.23~\mathrm{m/s}$. Without the action inspector, speed drops and collisions rise to $11\%$. Removing MPC increases speed to $22.88~\mathrm{m/s}$, but the collision rate remains high at $9\%$. These results highlight the full K-DQN’s strong balance of safety and efficiency.
@@ VI-B Normal Mode Validation
This section presents the experiments in normal mode (with seven initial vehicles in the roundabout in Fig. 5(a)) with comparison to benchmark DRL algorithms, PPO [23], A2C [37], ACKTR [38], and DQN [28].
Figure 10: Performance comparison with benchmarks in hard mode: (a) reward, (b) speed, and (c) collision rate.
Validation 1: Training performance. To assess training performance, we test K-DQN and benchmarks using three random seeds and varied scenarios. Figure 8(a) shows K-DQN achieves the highest rewards (over 200) and fastest convergence, outperforming PPO ($\sim 125$), A2C and ACKTR (similar), and DQN ($\sim 80$). Figure 8(b) shows K-DQN maintains the highest and most stable speed ($22~\mathrm{m/s}$), while PPO has the lowest and most unstable. A2C reaches  $18~\mathrm{m/s}$; others are similar. Figure 8(c) shows K-DQN achieves the lowest collision rate (<0.05<0.05), compared to PPO and A2C ($\sim 0.2$), ACKTR ($\sim 0.35$), and DQN ($\sim 0.55$).
Validation 2: Reward values among evaluation. Figure 9 illustrates the reward comparison between the proposed K-DQN and other benchmark algorithms. DQN has the lowest reward during the evaluation, falling below 75. A2C and ACKTR are similar, both increasing the reward to around 75. PPO has a relatively higher reward of around 100, peaking at 125. The reward of the proposed K-DQN fluctuates around 175, significantly surpassing other benchmark algorithms.